Solve the equation |y-6|+2y=9
So far, I have:
y-6=-(9-2y)
y-6=-9+2y
-y=-3
y=3
I tested that, and it is correct. Am I on the right track, and/or is there another answer? How do I find this additional answer?
Also - Graph the equation y=|x+4| + |3-x|
I know how to graph, but I am not quite sure how to start off. Many thanks if you can simplify it.

Answer by stanbon(48546)

Solve the equation |y-6|+2y=9
Rearrange:
|y-6| = 9-2y
y-6 = 9-2y or y-6 = -(9-2y)= 2y-9
3y = 15 or y = 3
y = 5 or y = 3
----------------
Testing:
Does y = 5 ?
|5-6| + 2*5 = 9
1+10 = 9
False; so y=5 is not a solution.
------------------
Testing:
Does y = 3 ?
|3-6| + 2*3 = 9
3 + 6 = 9
True ; so y=3 is the only solution.
------------------
Cheers,
Stan H.